OpenStudy (anonymous):

(0.4x10^-6)(0.7x10^-2) simplify #medallife

sam (.sam.):

\[(0.4\times 10^{-6})(0.7\times 10^{-2}) \\ \\ \Large =(0.4 \times 0.7 )\times 10^{-6+(-2)}\]